The pressure of 2 moles of an ideal gas at 546k having volume 44.8 L is

Grade:

2 Answers

harsha vardhan
5 years ago
P=2atm..
W. K. T PV=nRT 
P=NRT/V...V=44.8L n=2 T=546K R=0.0821(since the temp is in Kelvin volume is in Litre ..i.e pressure will be in ATM. .)
P=2(0.0821)(546)/44.6
P=1092(0.0821)/44.6
P=24.375(0.0821)
P=2.0011 implies P= 2atm.. 
P=
Aastha Singh
5 years ago
dear student,

Apply gas law
PV=nRT
where, P= pressure
V= 44.8 Lt.
n= 2 moles
R= 0.0821 L atm mol-1k-1, gas constant
T= 546 k
P*44.8 = 2*0.0821*546
P= 2atm